# Degree-Weighted DeGroot Learning

### v2

@@ Abstract
We examine belief formation in large stochastic networks under a degree-weighted DeGroot learning process. In many social settings, agents may place more trust in well-connected individuals or, conversely, discount their influence. Existing research on random networks primarily assumes that agents assign equal weights to their neighbors’ opinions. We relax this assumption by allowing agents to weight neighbors based on their degree. Our main technical contribution is to derive the asymptotic properties of learning outcomes, particularly the consensus belief and convergence speed, under this degree-weighted framework. Using this result, we analyze how the weighting rule affects consensus beliefs, societal wisdom, and convergence speed. We find that a more popularity-favoring rule—i.e., assigning greater weight to higher-degree neighbors—harms wisdom but has a non-monotonic effect on convergence speed. Whether it accelerates or slows convergence depends on the diversity of views within high- and low-degree groups. This highlights a potential tradeoff between faster convergence and an unwise consensus in networks where agents favor highly connected neighbors.
@@ 1 Introduction
The DeGroot model of social learning in networks has been instrumental in studying societal belief dynamics, such as whether society reaches consensus (DeGroot, 1974), agents’ network position and their long-run influence on consensus (DeMarzo, Vayanos and Zwiebel, 2003), the “wisdom of the crowd” phenomenon (Acemoglu, Ozdaglar and ParandehGheibi, 2010; Golub and Jackson, 2010), and the speed of belief convergence (Golub and Jackson, 2012). In a typical DeGroot learning model, agents repeatedly aggregate the opinions of their neighbors according to fixed weights over time. To gain tractability—especially when examining learning dynamics in large networks—agents are often assumed to assign equal weight to the opinions of all their neighbors (Golub and Jackson, 2012). However, the equal-weighting DeGroot model, while a useful benchmark, lacks certain empirically relevant elements. For example, it cannot capture attitudes toward popularity, where society may overweight the opinions of highly connected agents, believing that these individuals are exposed to a larger volume of information. Conversely, equal weighting fails to capture sentiment resistant to popularity, where society may underweight the opinions of highly connected agents, perceiving them as having more redundant or less trustworthy information sources. [^1] In practice, platforms may intentionally downweight highly popular users to provide “less-prominent” users with more visibility. For example, Facebook’s News Feed algorithm prioritizes content from friends, family, and close connections over posts from popular pages or celebrities. Similarly, Amazon’s initiatives, such as Amazon Launchpad and A+ Content, support new sellers and startups by offering enhanced product descriptions and media, thereby increasing their exposure and credibility. This approach improves the chances for smaller sellers to compete against established brands. [^2] The association between degree and low weighting does not need to be causal for the theory to hold. A high degree alone does not necessarily justify assigning lower weight to someone’s influence. Instead, it may be that individuals with certain characteristics (e.g., public figures with broad but less specialized appeal) tend to accumulate a high number of in-links, leading others to assign lower weight to their opinions. We thank the anonymous referee for suggesting this point.
@@
This paper develops a degree-weighted DeGroot model in which agents assign weights to their neighbors’ opinions based on their degree. We allow the weight to increase with the degree of a neighbor, capturing the notion of popularity-favoring; conversely, a decreasing weight reflects the sentiment of popularity-resistance. [^3] When the weight is independent of degree, the model reduces to the classical equal-weighting case. Our main contribution is an asymptotic theorem that establishes that the learning outcomes (consensus belief and convergence speed) in our degree-weighted learning framework can be approximated using the “expected” learning matrix. [^4] Previous works, with Golub and Jackson (2012), Dasaratha (2020) and Mostagir and Siderius (2021) being the closest, do not address the convergence of highly non-linear terms arising from degree-dependent weighting. Building on this result, we extend the standard equal-weighting DeGroot learning framework to degree-weighted scenarios, introducing a new dimension of comparative statics and an additional degree of freedom in degree dependence, which was previously difficult to explore.
Key Takeaway. Our degree-weighted learning model reveals that degree dependence shapes both societal wisdom and the speed of belief convergence, highlighting a fundamental trade-off between the two. Stronger degree dependence amplifies the influence of high-degree individuals, potentially undermining wisdom and exhibiting a largely monotonic effect. Its impact on convergence speed, however, is non-monotonic: whether greater degree dependence accelerates convergence depends on the diversity of views within high-degree groups. These forces suggest a potential trade-off—greater degree dependence may lead to faster convergence, but at the cost of societal wisdom.
@@
Wisdom. By fully characterizing agents’ influence—i.e., how their initial opinions shape the consensus belief—and the consensus belief itself, we apply the necessary and sufficient condition for a society to be wise, defined as the consensus belief uncovering the true state, as established in Golub and Jackson (2010). We find that degree-dependent weighting can shift a society between wisdom and its absence.
@@
Convergence Speed. The degree dependence can have a highly non-monotonic impact on the speed at which societal beliefs reach consensus. We find that when agents exhibit a greater degree of popularity-favoring, society does not necessarily reach consensus faster. Whether increasing the degree of popularity-favoring accelerates or slows convergence depends on the diversity of views within high-degree and low-degree groups. When low-degree groups outnumber high-degree groups and different groups potentially hold different views, increasing the degree of popularity-favoring may accelerate convergence; the reverse occurs when high-degree groups outnumber low-degree groups.
Tradeoff Between Wisdom and Convergence Speed. These forces can create a tradeoff between faster convergence and a “wiser” society: while a more popularity-resistant society may foster greater wisdom, reaching it may take longer when many low-degree groups exist, each potentially holding different views. Similarly, a more popularity-favoring learning rule may accelerate belief convergence at the expense of wisdom.
@@
Our paper relates to a substantial body of literature on DeGroot learning in social networks. DeGroot (1974) was among the first to examine naïve learning, where agents repeatedly aggregate their neighbors’ opinions according to a fixed set of weights, and identified conditions for belief convergence to consensus. [^5] Early contributions include French Jr (1956) and Harary (1959). For a recent survey on social learning and a discussion of the distinction between DeGroot and Bayesian models, see Golub and Sadler (2016). Subsequent work has extended the DeGroot paradigm to incorporate time-varying weights Chatterjee and Seneta (1977), initial biases or prejudices over agents’ prior beliefs Friedkin and Johnsen (1990), and opinion aggregation restricted to neighbors whose beliefs lie within a certain distance from one’s own Hegselmann and Krause (2002), among other variations. DeMarzo, Vayanos and Zwiebel (2003) later revisited the DeGroot model in an economic context, grounding naïve learning in a quasi-Bayesian framework and highlighting the role of persuasion bias in belief aggregation, along with other important insights. Gao (2021) examined centrality-weighted opinion dynamics but did not analyze the impact of such degree dependence on wisdom and convergence speed. Moreover, while these studies assumed deterministic networks, we analyze the DeGroot model in large, stochastic networks.
Acemoglu, Ozdaglar and ParandehGheibi (2010) and Golub and Jackson (2010) studied belief convergence to the true state in stochastic networks. Acemoglu, Ozdaglar and ParandehGheibi (2010) analyzed how “forceful agents” and network structure influence the gap between the consensus belief and the true state, showing that this gap shrinks as beliefs converge faster. Golub and Jackson (2010) identified conditions for societal wisdom under DeGroot learning and found that wisdom is independent of convergence speed. Our framework complements this literature by demonstrating that a third factor—degree dependence—affects both wisdom and convergence speed, creating a tradeoff between the two.
Most relatedly, Golub and Jackson (2012) examined how network features affect convergence speed in large networks, finding that homophily slows belief convergence. They derived the large-network property of the second-largest eigenvalue in modulus (SLEM), which is inversely related to convergence speed, for equal-weighting learning matrices. We extend their analysis to degree-dependent weighting matrices.
@@
Cerreia-Vioglio, Corrao and Lanzani (2024) extended the analysis to nonlinear robust belief aggregators (as opposed to linear DeGroot learning) and identified conditions for convergence, consensus, and the wisdom of the crowd. They showed that society attains wisdom when the SLEM falls below a certain threshold. Their discussion of SLEM relied on the large-sample property result from Golub and Jackson (2012) and thus depended on the equal-weighting assumption.
@@
Summary. Our key technical contribution is deriving the large-sample properties of the consensus belief and SLEM for degree-weighted learning matrices, where the closest related work is Golub and Jackson (2012), which assumed equal weighting. This approach is not only of theoretical interest but also of economic relevance, allowing us to extend existing theories on consensus, the wisdom of the crowd, and convergence speed to settings with degree-dependent weighting. The analysis provides an additional degree of freedom for comparative statics on degree dependence and reveals a potential tradeoff between convergence speed and societal wisdom.
@@ 2 Model: Networks and Degree-Weighted Learning
[…]
DeGroot Learning Model. In the classical DeGroot (1974) learning model, an agent’s belief such as the probability of global warming or the safety of a COVID vaccine, is a simple linear aggregation of her neighbors’ beliefs in the last period. In other words, the belief vector $\mathbf{b}(t)$ at time $t\in\mathbb{N}=\{0,1,\ldots\}$ is given by:
@@
in which $\mathbf{b}_{0}\in\mathbb{R}^{n}$ with $\|\mathbf{b}_{0}\|_{2}=1$ is an initial belief vector and $\mathbf{T}$, the learning matrix, is an $n\times n$ row-stochastic matrix, with each $T_{ij}$ representing the weight that agent $i$ puts on the belief of agent $j$ [^7] The constraint $\|\mathbf{b}_{0}\|_{2}=1$ is a standardization. Any initial belief $\mathbf{b}_{0}$ can be expressed as $\mathbf{b}_{0}=\|\mathbf{b}_{0}\|_{2}\cdot\mathbf{b}_{0}/\|\mathbf{b}_{0}\|_{2}$. Therefore, this effectively allows for any choice of initial belief.
Equal-Weighting Benchmark. Typically, agents are assumed to assign equal weight to their neighbors (e.g., Golub and Jackson, 2010, 2012). Under this assumption, the learning matrix $\mathbf{T}$ takes the following form:
$T_{ij}=\frac{A_{ij}}{d_{i}(\mathbf{A})}.$
Degree-Weighted Updating. While equal-weighting serves as a convenient and tractable benchmark, it falls short in capturing some real-life patterns. For instance, people may assign greater weight to celebrity opinions, or conversely, discount the views of more popular neighbors for various reasons discussed earlier. These observations suggest a learning heuristic in which the updating process depends on the popularity, or degree, of each source.
@@
$T_{ij}=\frac{A_{ij}d_{j}^{\alpha}}{\sum\limits_{j^{\prime}}A_{ij^{\prime}}d_{j^{\prime}}^{\alpha}}\,,$ (1)
where $d_{j}=d_{j}(\mathbf{A})$ is agent $j$’s degree in $\mathbf{A}$ and $\alpha\in\mathbb{R}$ captures how a neighbor’s weight depends on their degree. [^8] We note two points here. First, we allow for more general functional forms of degree dependence, $\phi(\alpha,d)\colon\mathbb{R}^{2}\rightarrow\mathbb{R}$, as long as they satisfy certain regularity conditions. We refer readers to Appendix B.1 for a more detailed discussion. Second, we acknowledge empirical evidence (e.g., Breza, Chandrasekhar and Tahbaz-Salehi, 2018) suggesting that agents may have limited knowledge of network structures and, in particular, may not know their neighbors’ degrees exactly. Our theoretical results can, to some extent, accommodate this lack of knowledge. As we show in the Appendix, our theory does not require exact knowledge of a neighbor’s degree. Instead, it suffices that each person has an impression of the number of friends that other individuals have. More precisely, what is required is a nonlinear transformation of the degree, i.e., $\phi(\alpha,d_{i})$. While the $i$-th individual may not directly observe $d_{j}$, we assume that they can observe $\phi(\alpha,d_{j})$ based on their impression of the $j$-th individual. And we have the following cases:
  • $\alpha=0$ (equal-weighting): the benchmark case studied in the literature (e.g., Golub and Jackson, 2012);
@@
Random Networks: Stochastic Block Model. Although DeGroot learning with deterministic networks has been extensively studied in the literature, its properties under random networks are less well understood. Random networks, however, offer distinct appeal and advantages over deterministic ones. [^9] For example, researchers often face data limitations and rely on statistical models of network formation, where links are probabilistically determined (see Dasaratha, 2020 for more examples). Moreover, in real-world settings, agents may encounter uncertainty regarding whom they will be matched with, which is better modeled using random networks (as in Acemoglu, Ozdaglar and ParandehGheibi, 2010). We use the stochastic block model to generate random networks in our paper (see Holland, Laskey and Leinhardt, 1983 for a classical discussion and Lee and Wilkinson, 2019 for a recent survey on its development).
@@
Notes: A network generated by a stochastic block model with $3$ groups, $30$ agents in each. The linking probability between agents $i$ and $j$ is given by $p_{ij}=0.3$ if $i$ and $j$ belong to the same group, and $0.04$ otherwise.
Suppose there are $n$ agents divided into $m$ groups. Let $\mathbf{n}=(n_{1},\ldots,n_{m})$, with $n_{k}$ being the number of agents in group $k\in\{1,\ldots,m\}$. Let P be an $m\times m$ symmetric matrix, which takes the following form:
$\textbf{P}=\begin{bmatrix}p_{11}&p_{12}&\ldots&p_{1m}\\ p_{21}&p_{22}&\ldots&p_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ p_{m1}&p_{m2}&\ldots&p_{mm}\end{bmatrix}\,.$ (2)
P represents the meeting probabilities for each group, where $p_{ij}>0$ denotes the linking probability within the same group if $i=j$, or between different groups if $i\neq j$ [^10] We assume $p_{ij}>0$ in order for the expected adjacency matrix (definition follows) to be strongly connected and aperiodic to ensure convergence to consensus (see Proposition 1).
@@
Let $\mathbf{R}={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$ be the corresponding expected adjacency matrix. The expected degree of a node $i$ in some group $k_{i}$ is
@@
Together with any $\alpha$, we define the expected learning matrix, $\mathbf{T}^{*}(\mathbf{R},\alpha)$, [^11] We note here that the role of matrices $\mathbf{R}$ and $\mathbf{T}^{*}$ is purely instrumental. They do not represent the actual network or the updating process but instead provide a tractable approximation that facilitates the analysis of the expectation and limiting behavior of the opinion dynamics. such that
$T^{*}_{ij}=\frac{R_{ij}d_{j}^{\alpha}(\mathbf{R})}{\sum\limits_{j^{\prime}}R_{ij^{\prime}}d_{j^{\prime}}^{\alpha}(\mathbf{R})}\,\forall i,j.$ (3)
@@
Lastly, we impose the following assumptions to establish our asymptotic results:
Assumption 1 (Density).
Let $\tau_{n}=\min\limits_{i}\frac{d_{i}(\mathbf{R})}{n}$, we assume
$\lim_{n\rightarrow\infty}\frac{\tau_{n}}{\sqrt{\frac{\log n}{n}}}=\infty\,.$ (4)
Assumption 2 (Comparable Densities).
$\frac{\max_{1\leq i,j\leq m}\{p_{ij}\}}{\min_{1\leq i,j\leq m}\{p_{ij}\}}<\infty\,.$ (5)
Assumption 1 requires that the graph cannot be too sparse; hence, the minimal degree should be significantly larger than $\sqrt{n\log n}$. Consequently, this ensures that, with probability tending to 1, there exists a path connecting any pair of nodes. Assumption 2 requires that the nonzero connection probabilities in the network are comparable. While interaction probabilities can vary significantly, their ratios must remain bounded. We need both assumptions to guarantee beliefs converge (to consensus) in the long run as networks grow as well as that the realized networks are “close enough” to expected networks (our Theorem 1 in the next section). [^12] Our Assumption 2 aligns with Condition (4) in Definition 3 of Golub and Jackson (2012). Assumption 1 differs from the Condition (1) in Definition 3 of Golub and Jackson (2012), as it is specifically needed to prove the concentration of nonlinear terms, such as $d_{j}^{\alpha}$. We do not require Condition (2) of Definition 3 in Golub and Jackson (2012) as we work with non-zero linking probabilities $p_{ij}>0$.

### v3

@@ Abstract
We study social learning in which agents weight neighbors’ opinions differently based on their degrees, capturing situations in which agents place more trust in well-connected individuals or, conversely, discount their influence. We derive asymptotic properties of learning outcomes in large stochastic networks and analyze how the weighting rule affects societal wisdom and convergence speed. We find that assigning greater weight to higher-degree neighbors harms wisdom but has a non-monotonic effect on convergence speed, depending on the diversity of views within high- and low-degree groups, highlighting a potential trade-off between convergence speed and wisdom. (JEL D83, D85, Z12)
@@ 1 Introduction
The DeGroot model of social learning in networks has been instrumental in studying societal belief dynamics, such as whether society reaches consensus (DeGroot, 1974), agents’ network position and their long-run influence on consensus (DeMarzo, Vayanos and Zwiebel, 2003), the “wisdom of the crowd” phenomenon (Acemoglu, Ozdaglar and ParandehGheibi, 2010; Golub and Jackson, 2010), and the speed of belief convergence (Golub and Jackson, 2012). In a typical DeGroot learning model, agents repeatedly aggregate the opinions of their neighbors according to fixed weights over time. To gain tractability—especially when examining learning dynamics in large random networks—agents are often assumed to assign equal weight to the opinions of all their neighbors (Golub and Jackson, 2012). However, the equal-weighting DeGroot model, while a useful benchmark, lacks certain empirically relevant elements. For example, it cannot capture attitudes toward popularity, where society may overweight the opinions of highly connected agents, believing that these individuals are exposed to a larger volume of information. Conversely, equal weighting fails to capture sentiment resistant to popularity, where society may underweight the opinions of highly connected agents, perceiving them as having more redundant or less trustworthy information sources. [^1] In practice, platforms may intentionally downweight highly popular users to provide “less-prominent” users with more visibility. For example, Facebook’s News Feed algorithm prioritizes content from friends, family, and close connections over posts from popular pages or celebrities. Similarly, Amazon’s initiatives, such as Amazon Launchpad and A+ Content, support new sellers and startups by offering enhanced product descriptions and media, thereby increasing their exposure and credibility. This approach improves the chances for smaller sellers to compete against established brands. [^2] The association between degree and low weighting does not need to be causal for the theory to hold. A high degree alone does not necessarily justify assigning lower weight to someone’s influence. Instead, it may be that individuals with certain characteristics (e.g., public figures with broad but less specialized appeal) tend to accumulate a high number of in-links, leading others to assign lower weight to their opinions. We thank the anonymous referee for suggesting this point.
@@
This paper develops a degree-weighted DeGroot model in which agents assign weights to their neighbors’ opinions based on their degree. We allow the weight to increase with the degree of a neighbor, capturing the notion of popularity-favoring; conversely, a decreasing weight reflects the sentiment of popularity-resistance. [^3] In our setting, when the weight is independent of degree, the model reduces to the classical equal-weighting case. Our main contribution is an asymptotic theorem showing that, in our degree-weighted random learning framework, the realized learning outcomes—the consensus belief and the convergence speed—converge to those generated by the “expected” learning matrix. [^4] Previous works, among which Golub and Jackson (2012), Dasaratha (2020) and Mostagir and Siderius (2021) are the closest, do not address the convergence of highly nonlinear terms arising from degree-dependent weighting. Building on this result, we extend the standard equal-weighting DeGroot learning framework to degree-weighted scenarios, introducing a new dimension of comparative statics and an additional degree of freedom in degree dependence, which was previously difficult to explore.
Wisdom
@@
By fully characterizing agents’ influence—i.e., how their initial opinions shape the consensus belief—and the consensus belief itself, we apply the necessary and sufficient condition for a society to be wise, defined as the consensus belief uncovering the true state, as established in Golub and Jackson (2010). We find that degree-dependent weighting can shift a society between wisdom and its absence.
@@
Convergence Speed
The degree dependence can have a highly non-monotonic impact on the speed at which societal beliefs reach consensus. We find that when agents exhibit a greater degree of popularity-favoring, society does not necessarily reach consensus faster. Whether increasing the degree of popularity-favoring accelerates or slows convergence depends on the diversity of views within high-degree and low-degree groups. When the number of low-degree groups exceeds that of high-degree groups, and different groups may hold different views, increasing the degree of popularity-favoring may accelerate convergence; the reverse occurs when high-degree groups outnumber low-degree groups.
Trade-off Between Wisdom and Convergence Speed
These forces can create a trade-off between faster convergence and a “wiser” society: while a more popularity-resistant society may foster greater wisdom, reaching it can take longer when many low-degree groups exist, as each group potentially holds different views. Similarly, a more popularity-favoring learning rule may accelerate belief convergence at the expense of wisdom. The following table illustrates such a trade-off. [^5] We discuss the theoretical results underlying this table in more detail in Section 4.3.
Table 1: Trade-off between Wisdom and Convergence Speed.
# Groups Degree Dependence
Low High Popularity-resistant Popularity-favoring
($\alpha\rightarrow$-$\infty$) ($\alpha\rightarrow\infty$)
1 1 wise, fast unwise, fast
multiple 1 wise, slow unwise, fast
1 multiple wise, fast unwise, slow
Notes: The first two columns report the number of low- and high-degree groups. The parameter $\alpha$ governs the weight that agents place on high-degree neighbors: $\alpha\rightarrow$-$\infty$ corresponds to an extremely popularity-resistant learning rule, while $\alpha\rightarrow\infty$ corresponds to an extremely popularity-favoring rule. The entries summarize whether society converges to wisdom and whether convergence is fast or slow under the corresponding regime.
@@
Our paper relates to a substantial body of literature on DeGroot learning in social networks. DeGroot (1974) was among the first to examine naïve learning, where agents repeatedly aggregate their neighbors’ opinions according to a fixed set of weights, and identified conditions for belief convergence to consensus. [^6] Early contributions include French Jr (1956) and Harary (1959). For a recent survey on social learning and a discussion of the distinction between DeGroot and Bayesian models, see Golub and Sadler (2016). Subsequent work has extended the DeGroot paradigm to incorporate time-varying weights Chatterjee and Seneta (1977), initial biases or prejudices over agents’ prior beliefs Friedkin and Johnsen (1990), and opinion aggregation restricted to neighbors whose beliefs lie within a certain distance from one’s own Hegselmann and Krause (2002), among other variations. DeMarzo, Vayanos and Zwiebel (2003) later revisited the DeGroot model in an economic context, grounding naïve learning in a quasi-Bayesian framework and highlighting the role of persuasion bias in belief aggregation, along with other important insights. Gao (2021) examined centrality-weighted opinion dynamics but did not analyze the impact of such degree dependence on wisdom and convergence speed. While these studies assumed deterministic networks, we analyze the DeGroot model in large stochastic networks.
Acemoglu, Ozdaglar and ParandehGheibi (2010) and Golub and Jackson (2010) studied belief convergence to the true state in stochastic networks. Acemoglu, Ozdaglar and ParandehGheibi (2010) analyzed how “forceful agents” and network structure influence the gap between the consensus belief and the true state, showing that this gap shrinks as beliefs converge faster. Golub and Jackson (2010) identified conditions for societal wisdom under DeGroot learning and found logical independence between wisdom and convergence speed. Our framework complements this literature by demonstrating that a third factor—degree dependence—affects both wisdom and convergence speed, creating a trade-off between the two.
Most relatedly, Golub and Jackson (2012) examined how network features affect convergence speed in large random networks, finding that homophily slows belief convergence. They derived the large-network property of the second-largest eigenvalue in modulus (SLEM), which is inversely related to convergence speed, for equal-weighting learning matrices. We extend their analysis to degree-dependent weighting matrices.
@@
Cerreia-Vioglio, Corrao and Lanzani (2024) extended the analysis to a broad class of non-Bayesian belief-updating rules, with DeGroot learning as a special case, and characterized conditions for convergence, consensus, and the wisdom of crowds. They showed that society attained wisdom when the SLEM fell below a certain threshold. While their focus was on generalizing DeGroot learning and identifying conditions for wisdom in that setting, our analysis instead studies the large-sample behavior of the SLEM and the role of degree-dependent weighting.
@@ 2 Model: Networks and Degree-Weighted Learning
Consider a network with $n$ agents. The associated $n\times n$ adjacency matrix $\mathbf{A}$ is defined by: $A_{ij}=A_{ji}=1$, if there is a link between agents $i$ and $j$; otherwise, $A_{ij}=A_{ji}=0$. [^7] This suggests that the network is undirected. We discuss the implications of this assumption, as well as the directed case, in Section 5. Let $d_{i}(\mathbf{A})=\sum_{j}A_{ij}$ be the degree of agent $i$, i.e., the number of links (friends) $i$ has.
DeGroot Learning Model
In the classical DeGroot (1974) learning model, an agent’s belief such as the probability of global warming or the safety of a vaccine, is a simple linear aggregation of her neighbors’ beliefs in the last period. In other words, the belief vector $\mathbf{b}(t)$ at time $t\in\mathbb{N}=\{0,1,\ldots\}$ is given by:
@@
in which $\mathbf{b}_{0}\in\mathbb{R}^{n}$ is an initial belief vector and $\mathbf{T}$, the learning matrix, is an $n\times n$ row-stochastic matrix, with $T_{ij}$ representing the weight that agent $i$ puts on the belief of agent $j$.
Equal-Weighting Benchmark
Typically, agents are assumed to assign equal weight to their neighbors (e.g., Golub and Jackson, 2010, 2012). Under this assumption, the learning matrix $\mathbf{T}$ takes the form $T_{ij}={A_{ij}}/{d_{i}(\mathbf{A})}.$
Degree-Weighted Updating
While equal-weighting provides a convenient and tractable benchmark, it falls short of capturing certain real-life patterns discussed in the introduction. These observations suggest a learning heuristic in which the updating process depends on the popularity, or degree, of each source.
@@
(1) $T_{ij}=\frac{A_{ij}d_{j}^{\alpha}}{\sum\limits_{j^{\prime}}A_{ij^{\prime}}d_{j^{\prime}}^{\alpha}}\,$,
where $d_{j}=d_{j}(\mathbf{A})$ is agent $j$’s degree in $\mathbf{A}$ and $\alpha\in\mathbb{R}$ captures how a neighbor’s weight depends on their degree. [^8] We note two points here. First, we allow for more general functional forms of degree dependence, $\phi(\alpha,d)\colon\mathbb{R}^{2}\rightarrow\mathbb{R}$, as long as they satisfy certain regularity conditions. We refer readers to Appendix 1 for a more detailed discussion. Second, we acknowledge empirical evidence (e.g., Breza, Chandrasekhar and Tahbaz-Salehi, 2018) suggesting that agents may have limited knowledge of network structures and, in particular, may not know their neighbors’ degrees exactly. We leave it to future work to explore the implications of this type of uncertainty on social learning. We have the following cases:
  • $\alpha=0$ (equal-weighting): the benchmark case studied in the literature;
@@
Random Networks: Stochastic Block Model
Although DeGroot learning with deterministic networks has been extensively studied in the literature, its properties under random networks are less well understood. Random networks, however, offer distinct appeal and advantages over deterministic ones. [^9] For example, researchers often face data limitations and rely on statistical models of network formation, where links are probabilistically determined (see Dasaratha, 2020 for more examples). Moreover, in real-world settings, agents may encounter uncertainty regarding whom they will be matched with, which is better modeled using random networks (as in Acemoglu, Ozdaglar and ParandehGheibi, 2010). We use the stochastic block model to generate random networks in our paper. [^10] See Holland, Laskey and Leinhardt (1983) for a classical discussion and Lee and Wilkinson (2019) for a recent survey on its development.
@@
{figurenotes}
A network generated by a stochastic block model with $3$ groups, $30$ agents in each. The linking probability between agents $i$ and $j$ is given by $p_{ij}=0.3$ if $i$ and $j$ belong to the same group, and $0.04$ otherwise.
Suppose there are $n$ agents divided into $m$ groups. Let $\mathbf{n}=(n_{1},\ldots,n_{m})$, with $n_{k}$ being the number of agents in group $k\in\{1,\ldots,m\}$. Let P be an $m\times m$ symmetric matrix
(2) $\textbf{P}=\begin{bmatrix}p_{11}&p_{12}&\ldots&p_{1m}\\ p_{21}&p_{22}&\ldots&p_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ p_{m1}&p_{m2}&\ldots&p_{mm}\end{bmatrix}\,$,
where $p_{ij}\geq 0$ represents the linking probability between any agent in group $i$ and any agent in group $j$.
@@
Let $\mathbf{R}=(R_{ij})={\rm I}\kern-1.79993pt{\rm E}\mathbf{A}$ be the corresponding expected adjacency matrix. The expected degree of a node $i$ in some group $k_{i}$ is
@@
Together with any $\alpha$, we define the expected learning matrix, $\mathbf{T}^{*}(\mathbf{R},\alpha)$, [^11] We note here that the role of matrices $\mathbf{R}$ and $\mathbf{T}^{*}$ is purely instrumental. They do not represent the actual network or the updating process but instead provide a tractable approximation that facilitates the analysis of the expectation and limiting behavior of the opinion dynamics.
(3) $T^{*}_{ij}=\frac{R_{ij}d_{j}^{\alpha}(\mathbf{R})}{\sum\limits_{j^{\prime}}R_{ij^{\prime}}d_{j^{\prime}}^{\alpha}(\mathbf{R})}\,,\forall i,j.$
We adopt standard definitions of convergence, consensus, and convergence speed as follows:
Definition 1 (Belief Convergence and Consensus)
Beliefs are said to converge if, for any initial belief profile $b(0)$, the sequence $b(t)$ satisfies $\lim_{t\rightarrow\infty}b_{i}(t)$ exists for all $i$. The network reaches consensus if beliefs converge to a common value, that is, $\lim_{t\rightarrow\infty}b_{i}(t)=b^{*}$ for all $i$, where $b^{*}$ is a constant independent of $i$.
Definition 2 (Convergence Speed)
(4) $\max\limits_{\|\mathbf{b}_{0}\|_{\mathbf{s}}=1}\|(\mathbf{T}^{t}-\mathbf{T}^{\infty})\mathbf{b}_{0}\|_{\mathbf{s}},$
in which $\|\mathbf{v}\|_{\mathbf{s}}=\sqrt{\sum_{i=1}^{n}s_{i}\mathbf{v}_{i}^{2}},$ where $\mathbf{v}_{i}$ is the $i$-th entry of $\mathbf{v}$ and $s_{i}$ is agent $i$’s influence to be defined in Equation (5).
This expression quantifies the largest possible deviation between beliefs at time $t$ and the limiting belief, across all initial beliefs whose $\|\cdot\|_{\mathbf{s}}$ norm is standardized to 1.121212The constraint $\|\mathbf{b}_{0}\|_{\mathbf{s}}=1$ is a standardization. Any initial belief $\mathbf{b}_{0}$ can be expressed as $\mathbf{b}_{0}=\|\mathbf{b}_{0}\|_{\mathbf{s}}\cdot\mathbf{b}_{0}/\|\mathbf{b}_{0}\|_{\mathbf{s}}$. Therefore, this effectively allows for any choice of initial belief. A larger value indicates slower convergence. [^13] The use of the $\|\cdot\|_{\mathbf{s}}$ norm is inspired by a similar norm in Golub and Jackson (2012). It places greater weight on discrepancies in the positions of more influential agents relative to the limit. We adopt the $\|\cdot\|_{\mathbf{s}}$ norm because, in the $\mathbf{s}$}-weighted $L^{2}$ space, the operator $\mathbf{T}$ is self-adjoint and its eigenvectors are orthogonal, allowing us to apply standard techniques as in Diaconis and Stroock (1991) to prove the second part of Proposition 1.
Definition 3
A network is strongly connected if there exists a path connecting any two nodes of the network. A network is aperiodic if there is no integer $k>1$ dividing the length of every cycle of the network. A nonnegative square matrix $\mathbf{M}$ is primitive if there exists a positive integer $k$ such that $\mathbf{M}^{k}$ is strictly positive (i.e., all entries of $\mathbf{M}^{k}$ are greater than zero).
It is a standard result that a strongly connected network is primitive if and only if it is aperiodic (see, e.g., Theorem 1.4 in Seneta, 2006).
@@
Lastly, we impose the following assumptions to establish our asymptotic results. We let the number of groups, $m$, be fixed as $n\rightarrow\infty$.
Assumption 1 (Density)
$\lim_{n\rightarrow\infty}{\tau_{n}}/{\sqrt{{(\log n)}/{n}}}=\infty\,$, where $\tau_{n}=\min_{i}{d_{i}(\mathbf{R})}/{n}$.
Assumption 2 (Comparable degrees)
${\max_{1\leq i,j\leq m}\{p_{ij}\}}/{\tau_{n}}<\infty\,.$
Assumption 3 (Bounded spectral gap)
$\limsup_{n\rightarrow\infty}|\lambda_{2}(\mathbf{T}^{*})|<1\,.$
Assumption 1 requires that the graph is not too sparse; in particular, the minimal degree must be significantly larger than $\sqrt{n\log n}$. Assumption 2 requires that the degrees in the network are comparable. [^14] Assumption 2 aligns with Condition (4) in Definition 3 of Golub and Jackson (2012). Assumption 1 differs from Condition (1) in Definition 3 of Golub and Jackson (2012), as it is specifically needed to prove concentration of nonlinear terms, such as $d_{j}^{\alpha}$. Assumption 3 is similar to Condition (3) in Definition 3 of Golub and Jackson (2012), ensuring that $\mathbf{A}$ is primitive with high probability and that $\mathbf{R}$ is primitive for sufficiently large nn. [^15] See Appendix A for a formal statement and proof. Because $\lambda_{2}$ determines convergence speed (see Proposition 1), this assumption prevents convergence from becoming arbitrarily slow as the network grows. It is stated as a condition on eigenvalues; we provide alternative sufficient conditions based solely on model primitives in Online Appendix O4. All three assumptions ensure that, as networks grow, our characterization of learning outcomes in the deterministic case (as captured by Proposition 1) holds with high probability, and that learning outcomes in realized networks are sufficiently close to those in expected networks (Theorem 1).